Need Help with this Laplace transform

Solve IVP by the Laplace Transform: y+y=e2t , given y(0) = 0, y(0) = 1. a) Identify Y(s) = L{y}. 3) Solve for y(t).
Answer #1

y 10, 20, yo121 y ty- ezt = L( 8 +Y) = Llezt) s-2 ² Lly) - 1 +LIO) :) (5²+1) Lly) = + 1 s-2 1) LIJ) - + S (S-2) (52+1) 3

I use property of laplas transformation to solve this problem